# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ Appendix F Auxiliary results
Combining the bounds (F.1), (F.1), and (F.1) with the inequality (F.1), we obtain that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$ $\lesssim\Lambda\varepsilon d+\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$
$\quad+\left(1\vee\frac{L}{b}\right)\frac{(M+d)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}$ (42)
@@
is carried out in a similar way. As before, let us fix an arbitrary $\theta\in\Theta$, let $\theta_{\varepsilon}$ be the closest to $\theta$ element of $\Theta_{\varepsilon}$ and denote the log-potentials corresponding to $\theta$ and $\theta_{\varepsilon}$ by $\varphi$ and $\varphi_{\varepsilon}$, respectively. According to Lemma F.2, it holds that
@@
${\mathcal{T}}_{\infty}|\varphi-\varphi_{\varepsilon}|=\mathbb{E}_{Y\sim{\mathcal{N}}(m,\Sigma)}\left|\varphi(Y)-\varphi_{\varepsilon}(Y)\right|\leqslant\mathbb{E}_{Y\sim{\mathcal{N}}(m,\Sigma)}\left[\Lambda\varepsilon\big(1+\|Y\|^{2}\big)\right]=\Lambda\varepsilon(1+\|m\|^{2}+\mathrm{Tr}(\Sigma))\lesssim\Lambda\varepsilon.$
@@
${\mathcal{A}}(x,T)\leqslant\overline{{\mathcal{A}}}(T)=\left(\frac{be^{2}R^{2}}{\sqrt{d}}+4e^{2}\sqrt{d}\right)\arcsin(e^{-bT})-10e^{2}\sqrt{d}\log\big(1-e^{-2bT}\big)\quad\text{for all $x\in\mathrm{supp}(\rho_{0})$.}$
Note that the expression in the right-hand side is of order
@@
$\left|\log{\mathcal{T}}_{T}\big[e^{\varphi}\big](Z_{j})-\log{\mathcal{T}}_{T}\big[e^{\varphi_{\varepsilon}}\big](Z_{j})\right|\lesssim{\mathcal{H}}(T)\left(\Lambda\varepsilon\right)^{1/{\mathcal{K}}(T)}\quad\text{almost surely for all $j\in\{1,\dots,n\}$.}$ (43)
This yields that
@@
$\leqslant 2\mathrm{Var}\left(\log{\mathcal{T}}_{T}\big[e^{\varphi}\big](Z_{1})-\log{\mathcal{T}}_{T}\big[e^{\varphi^{*}}\big](Z_{1})\right)+2\mathrm{Var}\left(\log{\mathcal{T}}_{T}\big[e^{\varphi_{\varepsilon}}\big](Z_{1})-\log{\mathcal{T}}_{T}\big[e^{\varphi}\big](Z_{1})\right)$ (45)
@@
The inequalities (43), (F.1), and (F.1) combined with (F.1) yield that
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$
$\quad+{\mathcal{H}}(T)\left(\Lambda\varepsilon\right)^{1/{\mathcal{K}}(T)}+\frac{(M+1)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}$ (46)
@@
Step 4: choice of $\varepsilon$ and final bounds.  It remains to choose a proper $\varepsilon>0$ to finish the proof. The inequalities (F.1), (F.1), and $|\Phi_{\varepsilon}|\leqslant(2/\varepsilon)^{D}$ imply that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$ $\lesssim\Lambda\varepsilon d+\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(1/\varepsilon)+\log(1/\delta)\big)}{n}}$
$\quad+\left(1\vee\frac{L}{b}\right)\frac{(M+d)\big(D\log(1/\varepsilon)+\log(1/\delta)\big)\log n}{n}$ (47)
@@
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(1/\varepsilon)+\log(1/\delta)\big)}{n}}$
$\quad+{\mathcal{H}}(T)\left(\Lambda\varepsilon\right)^{1/{\mathcal{K}}(T)}+\frac{(M+1)\big(D\log(1/\varepsilon)+\log(1/\delta)\big)\log n}{n}$ (48)
@@
$\lesssim\log\Lambda+\log(nd)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)+M{\mathcal{K}}(T)+\left(1-\frac{1}{{\mathcal{K}}(T)}\right)\log(M+LR^{2}+Ld^{2})$
@@
Substituting this bound into (F.1) and (F.1), we obtain that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$ $\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
An upper bound on $|{\mathrm{d}}F(s)/{\mathrm{d}}s|$ will follow from Lemma F.10. Indeed, due to the conditions of the lemma, for any $y\in\mathbb{R}^{d}$, it holds that
@@
Applying Lemma F.10, we obtain that
@@
$\left|\frac{{\mathrm{d}}F(s)}{{\mathrm{d}}s}\right|\leqslant e^{{\mathcal{A}}(x,T)({\mathcal{K}}(T)+1/{\mathcal{K}}(T))+M{\mathcal{K}}(T)}\left(M+\frac{LR^{2}}{2}+\frac{L(50d+d^{2}/8)}{b}\right)^{1-1/{\mathcal{K}}(T)}\left({\mathcal{T}}_{\infty}|f_{1}-f_{0}|\right)^{1/{\mathcal{K}}(T)}$
@@
Applying Lemma F.4, we obtain that
@@
$\leqslant\mathbb{E}_{\eta}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+2e^{-bT}\sqrt{\lambda L}\,\eta^{\top}\Sigma^{-1/2}(x_{0}-m)+\frac{\lambda L(1-e^{-2bT})\|\eta\|^{2}}{b}\right\}$
$=e^{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)}\prod\limits_{j=1}^{d}\left(\mathbb{E}_{\eta_{j}}\exp\left\{2e^{-bT}\sqrt{\lambda L}\,\eta_{j}\left(\Sigma^{-1/2}(x_{0}-m)\right)_{j}+\frac{\lambda L(1-e^{-2bT})\eta_{j}^{2}}{b}\right\}\right).$
@@
$\leqslant\left(1-\frac{\lambda L(1-e^{-2bT})}{b}\right)^{-d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+\frac{2e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}}{1-\lambda L(1-e^{-2bT})/b}\right\}$
$\leqslant\left(1-\frac{\lambda L}{b}\right)^{-d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+\frac{2e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}}{1-\lambda L/b}\right\}.$
@@
$\mathbb{E}\left[e^{2\lambda L\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right]\leqslant 2^{d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+4e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\},$
@@
According to Lemma F.10, we have
@@
$\log\mathbb{E}\left[e^{u^{\top}\Sigma^{-1/2}(X_{T}^{*}-m)}\,\big|\,X_{0}=x_{0}\right]\leqslant M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+e^{-bT}\,u^{\top}\Sigma^{-1/2}(x_{0}-m)+\frac{(1-e^{-2bT})\|u\|^{2}}{4b}.$
@@
It remains to consider $g^{\prime\prime\prime}(\lambda)$ to finish the proof. Taking into account (52) and applying (F.5), we obtain that
@@
$=\mathsf{P}_{\lambda}\big(\xi-\mathbb{E}\xi\big)^{3}-3\mathsf{P}_{\lambda}\big(\xi-\mathbb{E}\xi\big)^{2}\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)+3\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)\left(\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)\right)^{2}-\left(\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)\right)^{3}$
@@
Thus, the right-hand side of (F.5) simplifies to
@@
$\mathsf{P}_{\lambda}e^{|\xi-\mathbb{E}\xi|/t}=\frac{\mathbb{E}e^{|\xi-\mathbb{E}\xi|/t+\lambda(\xi-\mathbb{E}\xi)}}{\mathbb{E}e^{\lambda(\xi-\mathbb{E}\xi)}}\leqslant\mathbb{E}\exp\left\{\frac{|\xi-\mathbb{E}\xi|}{t}+\lambda(\xi-\mathbb{E}\xi)\right\}\leqslant\mathbb{E}\exp\left\{\left(\frac{1}{t}+\lambda\right)|\xi-\mathbb{E}\xi|\right\}.$
@@
Applying Lemma F.7, we conclude that
@@
Appendix G Further Details of Numerical Experiments
In this section, we elaborate on details of numerical experiments presented in Section 3. The section is organized as follows. In Appendix G.1, we discuss general implementation details. Appendix G.2 presents information about the metrics used and describes the algorithm performance. Appendix G.3 provides additional details about the 25 Gaussian mixture experiment. Appendix G.4 provides additional information about the I2I translation experiment in the ALAE latent space. Finally, Appendix G.5 presents all final hyperparameter values used for the experiments.
G.1 General Implementation Details

### v3

@@ Appendix E Auxiliary results
Combining the bounds (E.1), (E.1), and (E.1) with the inequality (E.1), we obtain that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\Lambda\varepsilon d+\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$ (50)
$\quad+\left(1\vee\frac{L}{b}\right)\frac{(M+d)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}$
@@
is carried out in a similar way. As before, let us fix an arbitrary $\theta\in\Theta$, let $\theta_{\varepsilon}$ be the closest to $\theta$ element of $\Theta_{\varepsilon}$ and denote the log-potentials corresponding to $\theta$ and $\theta_{\varepsilon}$ by $\varphi$ and $\varphi_{\varepsilon}$, respectively. According to Lemma E.2, it holds that
@@
${\mathcal{T}}_{\infty}|\varphi-\varphi_{\varepsilon}|$ $=\mathbb{E}_{Y\sim{\mathcal{N}}(m,\Sigma)}\left|\varphi(Y)-\varphi_{\varepsilon}(Y)\right|$
$\leqslant\mathbb{E}_{Y\sim{\mathcal{N}}(m,\Sigma)}\left[\Lambda\varepsilon\big(1+\|Y\|^{2}\big)\right]$
$=\Lambda\varepsilon(1+\|m\|^{2}+\mathrm{Tr}(\Sigma))\lesssim\Lambda\varepsilon.$
@@
${\mathcal{A}}(x,T)\leqslant\overline{{\mathcal{A}}}(T)=\left(\frac{be^{2}R^{2}}{\sqrt{d}}+4e^{2}\sqrt{d}\right)\arcsin(e^{-bT})-10e^{2}\sqrt{d}\log\big(1-e^{-2bT}\big)$
for all $x\in\mathrm{supp}(\rho_{0})$. Note that the expression in the right-hand side is of order
@@
$\left|\log{\mathcal{T}}_{T}\big[e^{\varphi}\big](Z_{j})-\log{\mathcal{T}}_{T}\big[e^{\varphi_{\varepsilon}}\big](Z_{j})\right|\lesssim{\mathcal{H}}(T)\left(\Lambda\varepsilon\right)^{1/{\mathcal{K}}(T)}$ (51)
almost surely for all $j\in\{1,\dots,n\}$. This yields that
@@
$\leqslant 2\mathrm{Var}\left(\log{\mathcal{T}}_{T}\big[e^{\varphi}\big](Z_{1})-\log{\mathcal{T}}_{T}\big[e^{\varphi^{*}}\big](Z_{1})\right)$
$\quad+2\mathrm{Var}\left(\log{\mathcal{T}}_{T}\big[e^{\varphi_{\varepsilon}}\big](Z_{1})-\log{\mathcal{T}}_{T}\big[e^{\varphi}\big](Z_{1})\right)$ (53)
@@
The inequalities (51), (E.1), and (E.1) combined with (E.1) yield that
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$ (54)
$\quad+{\mathcal{H}}(T)\left(\Lambda\varepsilon\right)^{1/{\mathcal{K}}(T)}+\frac{(M+1)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}$
@@
Step 4: choice of $\varepsilon$ and final bounds.  It remains to choose a proper $\varepsilon>0$ to finish the proof. The inequalities (E.1), (E.1), and $|\Phi_{\varepsilon}|\leqslant(2/\varepsilon)^{D}$ imply that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\Lambda\varepsilon d+\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(1/\varepsilon)+\log(1/\delta)\big)}{n}}$ (55)
$\quad+\left(1\vee\frac{L}{b}\right)\frac{(M+d)\big(D\log(1/\varepsilon)+\log(1/\delta)\big)\log n}{n}$
@@
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(1/\varepsilon)+\log(1/\delta)\big)}{n}}$ (56)
$\quad+{\mathcal{H}}(T)\left(\Lambda\varepsilon\right)^{1/{\mathcal{K}}(T)}+\frac{(M+1)\big(D\log(1/\varepsilon)+\log(1/\delta)\big)\log n}{n}$
@@
$\lesssim\log\Lambda+\log(nd)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)$
$\quad+M{\mathcal{K}}(T)+\left(1-\frac{1}{{\mathcal{K}}(T)}\right)\log(M+LR^{2}+Ld^{2})$
@@
Substituting this bound into (E.1) and (E.1), we obtain that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
An upper bound on $|{\mathrm{d}}F(s)/{\mathrm{d}}s|$ will follow from Lemma E.10. Indeed, due to the conditions of the lemma, for any $y\in\mathbb{R}^{d}$, it holds that
@@
Applying Lemma E.10, we obtain that
@@
$\left|\frac{{\mathrm{d}}F(s)}{{\mathrm{d}}s}\right|$ $\leqslant\exp\left\{{\mathcal{A}}(x,T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)+M{\mathcal{K}}(T)\right\}$
$\quad\cdot\left(M+\frac{LR^{2}}{2}+\frac{L(50d+d^{2}/8)}{b}\right)^{1-1/{\mathcal{K}}(T)}\left({\mathcal{T}}_{\infty}|f_{1}-f_{0}|\right)^{1/{\mathcal{K}}(T)}$
@@
Applying Lemma E.4, we obtain that
@@
$\leqslant\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\cdot\mathbb{E}_{\eta}\exp\left\{2e^{-bT}\sqrt{\lambda L}\,\eta^{\top}\Sigma^{-1/2}(x_{0}-m)+\frac{\lambda L(1-e^{-2bT})\|\eta\|^{2}}{b}\right\}$
$=\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\cdot\prod\limits_{j=1}^{d}\left(\mathbb{E}_{\eta_{j}}\exp\left\{2e^{-bT}\sqrt{\lambda L}\,\eta_{j}\left(\Sigma^{-1/2}(x_{0}-m)\right)_{j}+\frac{\lambda L(1-e^{-2bT})\eta_{j}^{2}}{b}\right\}\right).$
@@
$\leqslant\left(1-\frac{\lambda L(1-e^{-2bT})}{b}\right)^{-d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\cdot\exp\left\{\frac{2e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}}{1-\lambda L(1-e^{-2bT})/b}\right\}$
$\leqslant\left(1-\frac{\lambda L}{b}\right)^{-d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\cdot\exp\left\{\frac{2e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}}{1-\lambda L(1-e^{-2bT})/b}\right\}.$
@@
$\mathbb{E}\left[e^{2\lambda L\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right]$ $\leqslant 2^{d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\cdot\exp\left\{4e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
According to Lemma E.10, we have
@@
$\log\mathbb{E}\left[e^{u^{\top}\Sigma^{-1/2}(X_{T}^{*}-m)}\,\big|\,X_{0}=x_{0}\right]$ $\leqslant M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)$
$+e^{-bT}\,u^{\top}\Sigma^{-1/2}(x_{0}-m)+\frac{(1-e^{-2bT})\|u\|^{2}}{4b}.$
@@
It remains to consider $g^{\prime\prime\prime}(\lambda)$ to finish the proof. Taking into account (60) and applying (E.5), we obtain that
@@
$=\mathsf{P}_{\lambda}\big(\xi-\mathbb{E}\xi\big)^{3}-3\mathsf{P}_{\lambda}\big(\xi-\mathbb{E}\xi\big)^{2}\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)$
$\quad+3\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)\left(\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)\right)^{2}-\left(\mathsf{P}_{\lambda}(\xi-\mathbb{E}\xi)\right)^{3}$
@@
Thus, the right-hand side of (E.5) simplifies to
@@
$\mathsf{P}_{\lambda}e^{|\xi-\mathbb{E}\xi|/t}=\frac{\mathbb{E}e^{|\xi-\mathbb{E}\xi|/t+\lambda(\xi-\mathbb{E}\xi)}}{\mathbb{E}e^{\lambda(\xi-\mathbb{E}\xi)}}$ $\leqslant\mathbb{E}\exp\left\{\frac{|\xi-\mathbb{E}\xi|}{t}+\lambda(\xi-\mathbb{E}\xi)\right\}$
$\leqslant\mathbb{E}\exp\left\{\left(\frac{1}{t}+\lambda\right)|\xi-\mathbb{E}\xi|\right\}.$
@@
Applying Lemma E.7, we conclude that
@@
Appendix F Further Details of Numerical Experiments
In this section, we elaborate on details of numerical experiments presented in Section 4. The section is organized as follows. In Appendix F.1, we discuss general implementation details. Appendix F.2 presents information about the metrics used and describes the algorithm performance. Appendix F.3 provides additional details about the 25 Gaussian mixture experiment. Appendix F.5 provides additional information about the I2I translation experiment in the ALAE latent space. Finally, Appendix F.6 presents all final hyperparameter values used for the experiments.
F.1 General Implementation Details